A wooden rectangular hollow box has outer edges of 6 inches and 5 inches, and height of 8 inches. The uniform thickness of the box is 1 inch.
a. find the amount of material used;
b. how many wooden blocks each cube with 1 inch edge will fit into the box?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

Volume of wood used = external volume - internal volume
External volume
L*B*H (6,5,8)
= 6*5*8= 240 in^3
Internal volume
subtract2 from each outer dimensions ( one for each edge) with lid
4*3*6= 72 in^3
wood used = 240-72=168 in^3
internal volume =72 cu.in
72 cubes will fit



Answer by ikleyn(52803)

.

When the problem comes formulated in this way,  it may  (and should)  perplex a reader.

It is because its formulation is  INCOMPLETE.

To be complete,  the problem must say if the box is open  (without the lid)  or closed  (with the lid).

But,  it looks like,  this  Math composer doesn't bother with such details.


Therefore,  the problem looks like a lame horse.